# Quantum information with quantum-like bits

### v1

@@ Appendix A Transformation of the emergent states to the computational basis
In ?? we discussed how a QL representation of quantum gates can be straightforwardly defined once the emergent states are mapped onto the computational basis. In general, this transformation cannot be expressed in analytic form as soon as the subgraphs defining the resources are not exactly regular. However, we show here that the problem of determining the map can be rewritten as a matrix inversion problem, easily solvable with standard numeric algebra tools.
@@
where $X^{(q)}\in\mathbb{R}^{{N_{G}}}$ is diagonal, with components $X_{ij}^{(q)}=\delta_{ij}x_{j}^{(q)}$, and similarly for the other three blocks in ??. ?? can be rewritten as the linear problem
@@
where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$, and where we made use of the generalization of ?? to arbitrary values of ${N_{\mathrm{QL}}}$:
$\psi_{\pm}^{(q)}=\tfrac{1}{\sqrt{2}}\begin{pmatrix}a_{1}^{(q)}\\ a_{2}^{(q)}\end{pmatrix}.$ (A.3)
?? is expressed in matrix form as
@@ Appendix C Spectral properties of QL observables
    $\bm{\sigma}\mapsto n=\sum_{q=1}^{N_{\mathrm{QL}}}2^{{N_{\mathrm{QL}}}-q}\left(\frac{1-\sigma^{(q)}}{2}\right)+1.$ (C.3)
    Let us define by $V_{1}$ the set of vectors orthogonal to the eigenvector $f_{1}$ of $\mathcal{F}$ with eigenvalue $\lambda_{1}$. Note that $\mathcal{F}$ maps $V_{1}$ onto itself, i.e., $\mathcal{F}\psi\in V_{1}$ for $\psi\in V_{1}$. In fact,

### v4

@@ Appendix A Transformation of the emergent states to the computational basis
In ?? we discussed how a QL representation of quantum gates can be straightforwardly defined once the emergent states are mapped onto the computational basis. In general, this transformation cannot be expressed in analytic form as soon as the subgraphs defining the resources are not exactly regular. However, we show here that the problem of determining the map can be rewritten as a simple matrix inversion problem.
@@
where $X^{(q)}\in\mathbb{R}^{{N_{G}}}$ is diagonal, with components $X_{ij}^{(q)}=\delta_{ij}x_{j}^{(q)}$, and similarly for the other three blocks in ??. ???? can be rewritten as the linear problem
@@
where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$. ?? is expressed in matrix form as
@@ Appendix C Spectral properties of QL observables
    $\bm{\sigma}\mapsto n=\sum_{l=0}^{{N_{\mathrm{QL}}}-1}2^{l}\delta_{\sigma_{l},\uparrow}.$ (C.3)
    We denote by $V_{1}$ the set of vectors orthogonal to the eigenvector $f_{1}$ of $\mathcal{F}$ with eigenvalue $\lambda_{1}$. Note that $\mathcal{F}$ maps $V_{1}$ onto itself, i.e., $\mathcal{F}\psi\in V_{1}$ for $\psi\in V_{1}$. In fact,
@@
Appendix D Controlled-NOT gate
In this appendix, we analyze the action of the CNOT gate on a pair of QL bits, $p,q\in{1,\dots,{N_{\mathrm{QL}}}}$. Here, we designate $p$ as the control QL bit, whose state determines whether to flip the state of the target QL bit $q$. For simplicity, we assume $p<q$; the case $p>q$ can be derived analogously. The unitary map implementing the transformation is
$U_{\mathrm{CNOT}}^{(p,q)}$ $=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.1)
where we defined the projection operators
$\mathcal{P}_{i}^{(p)}=\tfrac{1}{2}\left[U_{\mathcal{I}}^{(p)}+(-1)^{i}U_{\mathrm{Z}}^{(p)}\right],$ (D.2)
with $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$ and $i\in\{0,1\}$. These are mutually orthogonal [$\mathcal{P}_{\sigma}^{(p)}\mathcal{P}_{\sigma^{\prime}}^{(p)}=\delta_{\sigma\sigma^{\prime}}\mathcal{P}_{\sigma^{\prime}}^{(p)}$] and complementary [$\mathcal{P}_{\downarrow}^{(p)}+\mathcal{P}_{\uparrow}^{(p)}=U_{\mathcal{I}}^{(p)}$]. The action of ?? on the total resource ?? is
$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\{\uparrow,\downarrow\}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$ (D.3a)
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-q)}+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)}$ (D.3b)
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\left[\mathcal{R}_{\uparrow\downarrow}^{(p)}+\mathcal{R}_{\downarrow\uparrow}^{(p)}\right]\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.3c)
where $\mathcal{R}^{(p)}_{\sigma\pi}=\mathcal{P}^{(p)}_{\sigma}\mathcal{R}^{(p)}\left[\mathcal{P}^{(p)}_{\pi}\right]^{\dagger}$. As introduced in ??, we can minimize the computational complexity of this gate by assessing to which extent the structure of ?? can be approximated by a Cartesian product. This property enables information through transformations applied exclusively to individual one-QL-bit resources. Let us therefore study separately all terms which would break the Cartesian product construction. Starting from the two terms in ??, we can study the relative error
$\frac{\left\lVert\mathcal{R}_{\mathrm{X}}^{(q)}-\mathcal{R}^{(q)}\right\rVert}{\left\lVert\mathcal{R}^{(q)}\right\rVert}=\sqrt{\frac{8l{N_{G}}}{2k{N_{G}}+2l{N_{G}}}}=\frac{2}{\sqrt{1+\frac{k}{2l}}}\simeq 1+\frac{2l}{k}.$ (D.4)
In ?? we introduced the Frobenius norm, defined by $\left\lVert A\right\rVert^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}|A_{ij}|^{2}$, for each $A\in M_{m\times n}(\mathbb{C})$. Also, we noticed that
$\mathcal{R}^{(q)}_{\mathrm{X}}=U_{\mathrm{X}}^{[q]}\mathcal{R}^{(q)}\left[U_{\mathrm{X}}^{[q]}\right]^{\dagger}=\begin{pmatrix}A_{1}^{(q)}&-C^{(q)}\\ -\left[C^{(q)}\right]^{\top}&A_{2}^{(q)}\end{pmatrix}.$ (D.5)
?? is minimized by the parameter choice defined in ??. Finally, let us consider the cross contribution in ??. This terms involves the difference
$\mathcal{R}^{(p)}_{\uparrow\downarrow}+\mathcal{R}^{(p)}_{\downarrow\uparrow}$ $=\frac{1}{2}(\mathcal{R}^{(p)}-\mathcal{R}^{(p)}_{Z})$
$=\frac{1}{2}\begin{pmatrix}A_{1}^{(p)}-A_{2}^{(p)}&C^{(p)}-\left[C^{(p)}\right]^{\top}\\ \left[C^{(p)}\right]^{\top}-C^{(p)}&A_{2}^{(p)}-A_{1}^{(p)}\end{pmatrix},$ (D.6)
where
$\mathcal{R}^{(q)}_{\mathrm{Z}}=\begin{pmatrix}A_{2}^{(q)}&\left[C^{(q)}\right]^{\top}\\ C^{(q)}&A_{1}^{(q)}\end{pmatrix}$ (D.7)
is calculated similarly to ??. Without loss of generality it is possible to choose here
$A_{1}^{(p)}$ $=A_{2}^{(p)},$ (D.8a)
$C^{(p)}$ $=\left[C^{(p)}\right]^{\top},\hskip 14.22636pt\forall\;p=1,\cdots,{N_{\mathrm{QL}}}.$ (D.8b)
In essence, we obtain
$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\mathcal{S}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
$\quad+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}+\mathcal{O}(l/k)$ (D.9)
provided the conditions in ?? hold.